If the Federal Reserve purchases $1 billion dollars of government securities, how much does M1 increase if the required reserve ratio is 50% and banks do not hold excess reserves?

A) $0.5 billion

B) $1 billion

C) $2 billion

D) $5 billion
Answer #1

Answer

Multiplier =1/reserve ratio =1/0.5=2

change in M1=multiplier *amount of purchase of government securities

=2*1

=$2 billion

The M1 increases by $2 billion.

option C